OpenStudy (anonymous):

Can someone please help me? I just need like an example on how to subtract polynomials.

OpenStudy (campbell_st):

collect like terms \[3x^2 - 2 - ( 5x^2 + 4x) = (3 - 5)x^2 -4x - 2\]

OpenStudy (anonymous):

wait okay so is the thing after the = sign the result?

OpenStudy (anonymous):

I mean I'm basically just confused with the subtraction/combing of like terms. If I have the term -5x^2 in one polynomial and the term -8x^2 in another, would I subtract them to get 3x^2?

OpenStudy (anonymous):

@campbell_st

OpenStudy (campbell_st):

ok... here are 2 degree 1 polynomials P(x) = 3x - 2 H(x) = x + 3 find P(x) - H(x) = 3x - 2 - (x + 3) distribute the negative = 3x - 2 - x - 3 collect like terms = (3 - 1)x - 2 - 3 = 2x - 5 the same process is repeated no matter the degree of the polynomial

OpenStudy (anonymous):

@campell_st THANK YOU SO MUCH SORRY FOR THE LATE REPLY